OpenStudy (anonymous):

what is the x intercept and the y intercept of : 2y-4x-16=0

OpenStudy (ganpat):

2y - 4x - 16 = 0 put x = 0 and find y... this is y intercept put y = 0 and find x... this is x intercept..

OpenStudy (ganpat):

for instance.. for y intercept.. put x = 0; 2y - x (0) - 16 = 0 calculate for y !!

OpenStudy (anonymous):

Calculate the Y-Intercept : Notice that when x = 0 the value of y is 8/1 so this line "cuts" the y axis at y= 8.00000 y-intercept = 8/1 = 8.00000

OpenStudy (anonymous):

Calculate the X-Intercept : When y = 0 the value of x is 4/-1 Our line therefore "cuts" the x axis at x=-4.00000 x-intercept = 8/-2 = 4/-1 = -4.00000

OpenStudy (anonymous):

Calculate the Slope : Slope is defined as the change in y divided by the change in x. We note that for x=0, the value of y is 8.000 and for x=2.000, the value of y is 12.000. So, for a change of 2.000 in x (The change in x is sometimes referred to as "RUN") we get a change of 12.000 - 8.000 = 4.000 in y. (The change in y is sometimes referred to as "RISE" and the Slope is m = RISE / RUN) Slope = 4.000/2.000 = 2.000

OpenStudy (anonymous):

Geometric figure: Straight Line Slope = 4.000/2.000 = 2.000 x-intercept = 8/-2 = 4/-1 = -4.00000 y-intercept = 8/1 = 8.00000

OpenStudy (anonymous):

Please no direct answers. @shereenkhan

OpenStudy (anonymous):

@Rainbow_dashie

OpenStudy (rainbow_dashie):

Okay I think they explained it enough.

OpenStudy (anonymous):

@kittiwitti1

OpenStudy (kittiwitti1):

You don't need me now @saeda xD